an aeroplane covers a distance of 120 km in 1 hour 20 minutes find the average speed of aeroplane?
Answer by jim_thompson5910(35256)

1 hr 20 min = 4/3 hrs

d = rt

120 = r*(4/3)

120/(4/3) = r

r = 120/(4/3)

r = 90

Average Speed: 90 km/hr